OpenStudy (anonymous):

Solve 3 sqrtx+4 close sqrt – 2 = 13

OpenStudy (anonymous):

is that \[3*\sqrt{x+4} -2 = 13\]?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

okay, so we get \[3\sqrt{x+4} = 15\]\[\sqrt{x+4} = 5\]\[(\sqrt{x+4})² = 5²\]\[x = 21\]

OpenStudy (kropot72):

\[\sqrt[3]{x+4}-2=13\] \[x+4=15^{3}\] \[x=15^{3}-4\] x=3371